OpenStudy (anonymous):

How do you write an equation for the line passing through points (3,1) and (-2,6).

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/algebra/linear_equation/images/slope_given_2points/slope-of-a-line-diagram.gif what do you think the SLOPE for those 2 points is?

OpenStudy (anonymous):

-1

OpenStudy (jdoe0001):

ok, now pick either of those sets of points, (x1, y1) or (x2,y2) and stick it in the point-slope formula http://www.scs.sk.ca/cyber/elem/learningcommunity/math/math30aa/curr_content/matha30rev1/lesson3-4/PONITSLOPEFORMULA.JPG

OpenStudy (jdoe0001):

say pick the 1st ordered pair, and use -1 as your 'm', slope

OpenStudy (anonymous):

I plugged in using point (3,1) and getting answer y= -1x+4...not sure if the -1x is correct.

OpenStudy (jdoe0001):

well, => y-(1) = -1 (x-3) \( \implies y-1=-x+3 \implies y=-x+3+1 \implies y=-x+4 \ OR \ {\bf y=4-x} \)

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

well, => y-(1) = -1 (x-3) $$ \implies y-1=-x+3 \implies y=-x+3+1 \implies y=-x+4 \ OR \ {\bf y=4-x} $$

OpenStudy (anonymous):

okay...see what I'm doing wrong! Thank you so much...can go test now!